# A Distributed Consensus Algorithm for Prioritizing Autonomous Vehicle Passing at Unsignalized Intersections under Mixed Traffic

### v1

@@ Abstract
sectionIntroduction Research and development on autonomous vehicles is currently progressing rapidly, with the goal of commercializing fully autonomous driving technology at Level 4 or higher as defined by the Society of Automotive Engineers (SAE) in the United States [1, 2]. Furthermore, technologies for Connected Autonomous Vehicles (CAVs) are becoming essential beyond stand-alone autonomous driving [3].
@@
I Related Works
I-A Intersection Entry Priority
@@
I-B Distributed Consensus in CAV Environments
@@
II Consensus Algorithm Design
@@
II-A Intersection Consensus Flow
@@
II-A1 Vehicle Count-Based Consensus
@@
II-A2 Consensus Cycle and Re-consensus
@@
II-B Voting-based Consensus Model
@@
II-B1 Candidate Voting Process
@@
II-B2 Leader Election Voting Process
@@
II-C Distributed Consensus Scenario
@@
II-C1 Initial State at Unsignalized Intersections
All vehicles entering an unsignalized intersection simultaneously start in the $Init\text{-}Candidate$ state unlike the original Raft algorithm where all processes start in the $Follower$ state. Each vehicle initializes its address and direction and first casts a vote for itself. Subsequently, it simultaneously sends vote requests to all other vehicles and performs the Candidate Voting Process as described in Section II-B.
@@ II-C1 Initial State at Unsignalized Intersections
II-C2 Candidate Voting Response and State Transition
@@
II-C3 Leader Election Voting Response and State Transition
@@
II-C4 Termination of the Consensus Cycle and Re-consensus
@@
II-D Analysis of Safety and Liveness
@@ II-D2 Guarantee of Liveness
II-D3 Optimal Quorum for Safety and Liveness Guarantee
@@
Therefore, a practical solution is to ensure that consensus can be achieved at least among the responsive CAVs within the quorum. In this study, we set the minimum quorum as $k>\left\lceil\frac{N}{2}\right\rceil$, allowing safe agreement and liveness among the normally responding vehicles. Setting this minimum quorum is essential to enable rapid consensus among responsive CAVs and prevent congestion at intersections. The effectiveness of this quorum threshold in intersections where HVs and unresponsive autonomous vehicles coexist is evaluated through simulation experiments in the following Section III.
III implementation and Evaluation

### v2

@@ Abstract
I Introduction
Research and development on autonomous vehicles is currently progressing rapidly, with the goal of commercializing fully autonomous driving technology at Level 4 or higher as defined by the Society of Automotive Engineers (SAE) in the United States [1, 2]. Furthermore, technologies for Connected Autonomous Vehicles (CAVs) are becoming essential beyond stand-alone autonomous driving [3].
@@ I Introduction
II Related Works
II-A Intersection Entry Priority
@@
II-B Distributed Consensus in CAV Environments
@@
III Consensus Algorithm Design
@@
III-A Intersection Consensus Flow
@@
III-A1 Vehicle Count-Based Consensus
@@
III-A2 Consensus Cycle and Re-consensus
@@
III-B Voting-based Consensus Model
@@
III-B1 Candidate Voting Process
@@
III-B2 Leader Election Voting Process
@@
III-C Distributed Consensus Scenario
@@
III-C1 Initial State at Unsignalized Intersections
All vehicles entering an unsignalized intersection simultaneously start in the $Init\text{-}Candidate$ state unlike the original Raft algorithm where all processes start in the $Follower$ state. Each vehicle initializes its address and direction and first casts a vote for itself. Subsequently, it simultaneously sends vote requests to all other vehicles and performs the Candidate Voting Process as described in Section III-B.
@@ III-C1 Initial State at Unsignalized Intersections
III-C2 Candidate Voting Response and State Transition
@@
III-C3 Leader Election Voting Response and State Transition
@@
III-C4 Termination of the Consensus Cycle and Re-consensus
@@
III-D Analysis of Safety and Liveness
@@ III-D2 Guarantee of Liveness
III-D3 Optimal Quorum for Safety and Liveness Guarantee
@@
Therefore, a practical solution is to ensure that consensus can be achieved at least among the responsive CAVs within the quorum. In this study, we set the minimum quorum as $k>\left\lceil\frac{N}{2}\right\rceil$, allowing safe agreement and liveness among the normally responding vehicles. Setting this minimum quorum is essential to enable rapid consensus among responsive CAVs and prevent congestion at intersections. The effectiveness of this quorum threshold in intersections where HVs and unresponsive autonomous vehicles coexist is evaluated through simulation experiments in the following Section IV.
IV implementation and Evaluation